### v1

# ShuttleSet22: Benchmarking Stroke Forecasting with Stroke-Level Badminton Dataset

Abstract
In recent years, badminton analytics has drawn attention due to the advancement of artificial intelligence and the efficiency of data collection. While there is a line of effective applications to improve and investigate player performance, there are only a few public badminton datasets that can be used for researchers outside the badminton domain. Existing badminton singles datasets focus on specific matchups; however, they cannot provide comprehensive studies on different players and various matchups. In this paper, we provide a badminton singles dataset, ShuttleSet22, which is collected from high-ranking matches in 2022. ShuttleSet22 consists of 30,172 strokes in 2,888 rallies in the training set, 1,400 strokes in 450 rallies in the validation set, and 2,040 strokes in 654 rallies in the testing set with detailed stroke-level metadata within a rally. To benchmark existing work with ShuttleSet22, we test the state-of-the-art stroke forecasting approach, ShuttleNet, with the corresponding stroke forecasting task, i.e., predict the future strokes based on the given strokes of each rally. We also hold a challenge, Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies, at CoachAI Badminton Challenge 2023 to boost researchers to tackle this problem. The baseline codes and the dataset will be made available at the GitHub repo [^1] https://github.com/wywyWang/CoachAI-Projects/tree/main/CoachAI-Challenge-IJCAI2023/Track%202%3A%20Stroke%20Forecasting.
1 Introduction
Sports analytics has captured increasing attention since the advancement of technology brings great convenience for data collection and increases data diversity. In recent years, there is a line of studies applying advanced artificial intelligence techniques, e.g., computer vision on video frames Kim et al. (2022), motion tracking with sensors Ito and Gotoh (2018); Pillitteri et al. (2021), and machine learning models for action valuing Decroos et al. (2019); Merhej et al. (2021). Badminton serves as one of the major racket sports in the world in terms of participation, and it demands high physical and tactical conditions, which has attracted researchers on introducing novel applications Wang et al. (2020, 2021). For instance, Wang et al. (2022a) proposes long short-term extractors to quantify the win influence of each shot within a rally. Chang et al. (2022) designs the movement forecasting task to predict the players’ movements with graph-based approaches.
In this paper, we aim to release, ShuttleSet22, a stroke-level badminton singles dataset collected from high-ranking matches in 2022. ShuttleSet22 is an extension from ShuttleSet Wang et al. (2023) encompassing with 30,172 strokes (2,888 rallies), 1,400 strokes (450 rallies) in the validation set, and 2,040 strokes (654 rallies) in the testing set. While ShuttleSet22 shares similar stroke-level data formats with ShuttleSet, ShuttleSet22 consists of matches in 2022 instead of between 2018 and 2021. Therefore, ShuttleSet can be viewed as the latest iteration of the badminton singles dataset that can be used to examine the model’s effectiveness in recent matches. ShuttleSet22 is obtained from the public videos [^2] http://bwf.tv/ and is labeled by domain experts with the shot-by-shot labeling tool Huang et al. (2022).
To boost researchers’ engagements in badminton analytics, we, for the first time, create a challenge in CoachAI Badminton Challenge 2023333https://sites.google.com/view/coachai-challenge-2023/ in conjunction with IJCAI 2023. Specifically, we hold the forecasting of future turn-based strokes in badminton rallies (Track 2) and provide the state-of-the-art stroke forecasting approach Wang et al. (2022b) as the official baseline. This track has attracted about 100 participants to join for improving the effectiveness of the stroke forecasting task.
The paper is organized as follows: Related work is described in Section 2, the details of the stroke forecasting task, the data exploration, and the official baseline are introduced in Section 3, methods of participants are summarized in Section 4, the results from participants are depicted in Section 5, and we conclude our findings in Section 6. We select it as the official baseline due to not only the state-of-the-art performance but also the first work for turn-based sequence prediction.
2 Related Work: Public Badminton Datasets
Table 1: Comparison of previous badminton datasets with our datasets.
# Players # Match # Rally # Stroke Year
ShuttleSet22 (Ours) 35 58 3,992 33,612 2022
BadmintonDB Ban et al. (2022) 2 9 811 9,671 2018-2020
Generally, there are only a few public badminton datasets due to the heavy cost of collecting and labeling fine-grained records by domain experts Wang (2022). Recently, researchers have released badminton datasets to foster the sports community. For instance, the shuttlecock datasets Cartron (2022) contains 8K images of shuttlecocks which are resized to 640*640 pixels. This dataset includes the position of shuttlecocks which could train an object detection model. The BadmintonDB Ban et al. (2022) features rally, strokes, and outcome annotations between two players, which can be used in player-specific match analysis and prediction tasks, which is also the stroke-level dataset. However, BadmintonDB only consists of the same matchup (i.e., Kento Momota and Anthony Sinisuka Ginting) instead of the various matchups. Moreover, BadmintonDB collects matches from 2018 to 2020, while our ShuttleSet22 collects high-ranking matches in 2022 to reflect the state-of-the-art tactic records. We summarize the discrepancies between our ShuttleSet and BadmintonDB in Table 1.
3 Task Details
Figure 1: Illustration of Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies.
The goal of this shared task is to forecasting future turn-based strokes in badminton rallies, which aims to design forecasting models that are able to predict future strokes including shot types and locations based on the past stroke sequences. Figure 1 shows the whole flowchart of our task. The task page is available on Codalab.
3.1 Problem Formulation
Following Wang et al. (2022b), for each singles rally, given the observed $\tau$ strokes with type-area pairs and two players, the goal is to predict the future strokes including shot types and area coordinates for the next $n$ steps. To simplify the problem, $\tau$ is set to be 4 in our challenge, and $n$ is various based on the length of the rally.
Let $R=\{S_{r},P_{r}\}_{r=1}^{|R|}$ denote historical rallies of badminton matches, where the $r$-th rally is composed of a stroke sequence with type-area pairs $S_{r}=(\langle s_{1},a_{1}\rangle,\cdots,\langle s_{|S_{r}|},a_{|S_{r}|}\rangle)$ and a player sequence $P_{r}=(p_{1},\cdots,p_{|S_{r}|})$. At the $i$-th stroke, $s_{i}$ represents the shot type, $a_{i}=\langle x_{i},y_{i}\rangle\in\mathbb{R}^{2}$ are the coordinates of the shuttle destinations, and $p_{i}$ is the player who hits the shuttle. We denote Player A as the served player and Player B as the other for each rally in this paper. For instance, given a singles rally between Player A and Player B, $P_{r}$ may become $(A,B,\cdots,A,B)$. We formulate the problem of stroke forecasting as follows. For each rally, given the observed $\tau$ strokes $(\langle s_{i},a_{i}\rangle)_{i=1}^{\tau}$ with players $(p_{i})_{i=1}^{\tau}$, the goal is to predict the future strokes including shot types and area coordinates for the next $n$ steps, i.e., $(\langle s_{i},a_{i}\rangle)_{i={\tau+1}}^{\tau+n}$. We note that $n$ is pre-defined as the actual length of the corresponding rally.
3.2 Exploratory Data Analysis
We had released 58 matches including about 4,000 rallies with shot-level records following the BLSR format Wang et al. (2022a) for this task, where parts of the matches from 2018 to 2021 are the same as Wang et al. (2022b), while we add new matches collected from 2022.
Table 2: Explanation and example of each column.
Column Name Explanation
Rally Serial number of rallies in a match.
Ball Round The order of the shot in a rally.
Time The shot’s hitting time, which is in the format of hr:min:sec.
Frame num Equals to sec multiply fps.
Round score A Player A’s current score in the set.
Round score B Player B’s current score in the set.
Player The player who performed the shot.
Type The type of shot.
Aroundhead Hit the shuttle around the head or not.
Backhand Hit the shuttle with the backhand or not.
Landing height If the shuttle destination is hit above (1) or below (0) the net.
Landing area The grid of the shuttle destinations
Landing x
Landing y
The coordinates of the shuttle destinations.
Reason for loss The reason why the rally ended.
Get the point player The player who won the rally.
Player location area The location of the player who performed the shot.
Player location x,
Player location y
The coordinates of the player’s location when the player hits the shuttle.
Opponent location area The location of the player who prepared to receive the shot.
Opponent location x, Opponent location y The coordinates of the opponent’s location when the player hits the shuttle.
Set The current set in a match (best-of-3).
Match id Serial number of matches
Rally id Serial number of all rallies
Rally length Number of shots in a rally, rally_length-4 is the shots needed to predict.
Table 2 summarizes the definition of each column, which is mainly followed by Wang et al. (2022a). Figure 2 shows the heatmap of player location and opponent location, which is obvious that most of the players stay in both centers of the court and four corners of the court when they are going to attack. On the contrary, most of the players stay in the center of the court when defending so that they can immediately react to any kind of shot.
Table 3 indicates the number of different strike orders by shot type. It is observed that all shot types, except Clear and Smash, do not hit around the head. More than half of Push/Rush, Defensive Shot, and Short Service is hit with the backhand, while other shot types are hit with the forehand. The service type (i.e., how the player starts the first shot) has the most distinct difference in hitting position, with almost all short services hitting below the net and almost all long services hitting above the net.
Figure 2: The heatmap of player location and opponent location in our datasets.
Table 3: Shot type statistics.
Shot Type Not Around Head Around Head Forehand Backhand Higher than the net Below than the net Total Count
Drive 891 41 701 231 497 435 932
Net Shot 5721 4 3110 2615 158 5566 5725
Lob 5205 2 2706 2501 4890 316 5207
Clear 2062 1016 2808 270 2868 210 3078
Drop 2539 643 2863 319 133 3049 3182
Push/Rush 1994 17 992 1019 1277 734 2011
Smash 2402 1280 3641 41 56 3625 3682
Defensive Shot 4079 8 1742 2345 940 3146 4087
Short Service 1620 0 146 1474 21 1599 1620
Long Service 648 0 528 120 644 3 648
3.3 Evaluation Metrics
The evaluation scenarios will be evaluated by cross-entropy for shot type prediction and mean absolute error (MAE) for area coordinates prediction similar to the original work. Since this task exists stochasticity, we will also ask each team to generate 6 predicted sequences of each rally and take the closest one to ground truth for evaluation. We note that the original work of the stroke forecasting task is to generate 10 sequences, which we change to 6 for efficiency.
3.4 Official Baseline
Generally, ShuttleNet Wang et al. (2022b) is a position-aware fusion of rally progress and player styles framework consisting of Transformer-based architectures. ShuttleNet separates the contexts of both players to consider the overall playing habits for predicting the next strokes.
4 Participating Systems
There are about 95 participants joining the challenge system, and 16 teams submitted their testing results for the final phase in the CoachAI Badminton Challenge 2023 (Track 2). Most teams modified the official baseline (i.e., ShuttleNet) to approach the task. The brief descriptions of the submitted methods are summarized in Table 4.
5 Results
Table 4: Performance and brief descriptions of the stroke forecasting task in CoachAI Badminton Challenge 2023 (Track 2).
$\tau=4$
Rank Team Descriptions CE MAE Total
1 Intro_to_AI_team8 Since the embedded player number appears in the original ShuttleNet structure multiple times, we doubt that we can remove some of them to make our model simpler. Similarly, the embedded area has gone through many sigmoid and relu functions repetitively before passing them to the neural network. We simply ignore all the unnecessary activation functions and embedding processes in the original structure. 1.7892 0.7884 2.5776
2 Badminseok The given input parameters include ”type,” ”aroundhead,” ”backhand,” ”landing height,” ”landing_x,” ”landing_y,” ”player_location_area,” and ”opponent_location_area.” The desired output parameters are the same as the input parameters. A total loss function is defined as a weighted combination of two components: ”loss_area” and ”loss_type.” The weights are controlled by the alpha value, which can be adjusted for tuning. Additional losses, such as ”loss_aroundhead,” ”loss_backhand,” ”loss_height,” ”loss_playerlocationarea,” and ”loss_opponentlocationarea,” are included. The final ideal loss is obtained by combining the weighted losses of the two models. Hyperparameter tuning is performed separately for model 1 (optimizing loss area) and model 2 (optimizing loss type). The ensemble model is created by combining the outputs of both models to generate the final data. Additionally, we introduced two forward layers that returns vectors of area and shot to deepen the encoder and the decoder depth. The modified model showed better performance on regression task. 1.8127 0.7703 2.5830
3 NYCU-group4 Optimizing the parameters of the original baseline code, and testing the accuracy of the model on validation sets to obtain the model with the optimal weights. 1.8411 0.7826 2.6237
4 YHY We basically use the ShuttleNet from the paper to predict the future steps of type and landing coordinate. 1.9685 0.6797 2.6482
5 20 Use 9 convolution in FFN and add relative positional bias. 1.9366 0.7490 2.6856
6 LOL data augmentation using spatial mirroring and finding matching technique, added player area as input feature, change of activation function from baseline code. 1.9390 0.7618 2.7008
7 Group27 Based on ShuttleNet, we add Convolutional Unit and use Bayesian Optimization to find the optimal hyperparameter. 2.0200 0.7139 2.7339
8 Team_13 Based on Shuttlenet and adjust parameters. 1.9681 0.7671 2.7352
9 LinDan Train on extra epochs. 2.0097 0.7743 2.7841
10 Intro_to_AI_group_5 We apply some feature extraction methods to integrate features provided in the dataset. In comparison to the original ShuttleNet model, our work generates predictions based on more features, including shot angle, time difference, etc. 1.9710 0.8726 2.8436
11 14 We use ShuttleNet as our main approach in this competition. 2.1718 0.7013 2.8731
- ShuttleNet Wang et al. (2022b) 2.1777 0.6997 2.8774
12 GD_Wang Discretize location for ball type attention & curriculum learning with multi-task (base on ShuttleNet). 2.2479 0.7081 2.9560
13 ShuttleFold We commence our methodology by detailing the dataset employed and the associated data preprocessing steps. Following this, drawing upon the resemblances with AlphaFold2, we elucidate the initialization of our unique player and rally representations. Subsequently, we delve into the intricacies of the Evoformer, which facilitates the update of these representations. 2.1277 1.0107 3.1385
14 Awesome Badminton Increase embeddings dimension. 2.6579 0.7120 3.3699
15 Badminton is all you need Autoregressive model with SOTA techniques that will reveal soon, only trained in under 200 epochs. 2.6579 0.7120 3.3699
16 Group_28 Select several significant features for the past 4 stroke sequence and train a model with the encoder-decoder architecture for a single future sequence prediction including landing coordinates. Use the selected- predicted features as attributes for NN categorical classification on type of strokes. We iterates over the two continuously until the ball round ends. 4.6615 1.0268 5.6883
Table 4 reports the official leaderboard of the participants’ methods. We can observe that Team Intro_to_AI_team 8 outperforms the competition slightly than Team Badminseok, while Team YHY performs the best in terms of area predictions. There are 11 teams that perform better than ShuttleNet; however, we notice that these methods are built on top of ShuttleNet, which demonstrates the flexibility of ShuttleNet and some designs (e.g., hyper-parameters and activation functions) can be improved. In addition, the improvements of the participating teams mainly focus on the shot type prediction (2.1777 $\rightarrow$ 1.7892); however, the performance of area coordinates only slightly performs better than ShuttleNet (0.6997 $\rightarrow$ 0.6797), where most teams are inferior to the baseline. This points out the challenge of effectively incorporating two predictions, which is a potential future direction to explore.
6 Conclusion
In this paper, we propose ShuttleSet22, an extended dataset from ShuttleSet with stroke-level badminton singles records. ShuttleSet22 consists of fine-grained metadata for researchers to utilize various aspects for discovering thoughtful insights. Exploratory data analysis is conducted to illustrate the fundamental compositions of the dataset to bridge the gap between researchers outside the badminton domain. To foster researchers to incorporate advanced techniques with badminton analytics, we hold Track 2, a challenge in CoachAI Badminton Challenge 2023 that focuses on improving the effectiveness of stroke forecasting. In addition, we replicate the state-of-the-art baseline for the task as the official baseline. Methods and insights from participants are summarized and discussed to provide potential improvements for future research.

### v3

@@
-ShuttleSet22: Benchmarking Stroke Forecasting with Stroke-Level Badminton Dataset
+Benchmarking Stroke Forecasting with Stroke-Level Badminton Dataset
 Abstract
 In recent years, badminton analytics has drawn attention due to the advancement of artificial intelligence and the efficiency of data collection.
-While there is a line of effective applications to improve and investigate player performance, there are only a few public badminton datasets that can be used for researchers outside the badminton domain.
+While there is a line of effective applications to improve and investigate player performance, there are only a few public badminton datasets that can be used by researchers outside the badminton domain.
 Existing badminton singles datasets focus on specific matchups; however, they cannot provide comprehensive studies on different players and various matchups.
 In this paper, we provide a badminton singles dataset, ShuttleSet22, which is collected from high-ranking matches in 2022.
-ShuttleSet22 consists of 30,172 strokes in 2,888 rallies in the training set, 1,400 strokes in 450 rallies in the validation set, and 2,040 strokes in 654 rallies in the testing set with detailed stroke-level metadata within a rally.
-To benchmark existing work with ShuttleSet22, we test the state-of-the-art stroke forecasting approach, ShuttleNet, with the corresponding stroke forecasting task, i.e., predict the future strokes based on the given strokes of each rally.
-We also hold a challenge, Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies, at CoachAI Badminton Challenge 2023 to boost researchers to tackle this problem.
-The baseline codes and the dataset will be made available at the GitHub repo [^1] https://github.com/wywyWang/CoachAI-Projects/tree/main/CoachAI-Challenge-IJCAI2023/Track%202%3A%20Stroke%20Forecasting.
+ShuttleSet22 consists of 30,172 strokes in 2,888 rallies in the training set, 1,400 strokes in 450 rallies in the validation set, and 2,040 strokes in 654 rallies in the testing set, with detailed stroke-level metadata within a rally.
+To benchmark existing work with ShuttleSet22, we hold a challenge, Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies, at CoachAI Badminton Challenge @ IJCAI 2023, to encourage researchers to tackle this real-world problem through innovative approaches and to summarize insights between the state-of-the-art baseline and improved techniques, exchanging inspiring ideas.
+The baseline codes and the dataset are made available at the GitHub repo [^1] https://github.com/wywyWang/CoachAI-Projects/tree/main/CoachAI-Challenge-IJCAI2023.
 1 Introduction
-Sports analytics has captured increasing attention since the advancement of technology brings great convenience for data collection and increases data diversity.
-In recent years, there is a line of studies applying advanced artificial intelligence techniques, e.g., computer vision on video frames Kim et al.
-(2022), motion tracking with sensors Ito and Gotoh (2018); Pillitteri et al.
+Sports analytics has garnered increasing attention since the advancement of technology, which greatly facilitates data collection and increases data diversity.
+In recent years, there has been a surge in studies applying advanced artificial intelligence techniques, e.g., computer vision on video frames Kim et al.
 (2022), and machine learning models for action valuing Decroos et al.
@@
 (2021).
-Badminton serves as one of the major racket sports in the world in terms of participation, and it demands high physical and tactical conditions, which has attracted researchers on introducing novel applications Wang et al.
-(2020, 2021).
+Badminton, as one of the major racket sports worldwide in terms of participation, demands high physical and tactical conditions, attracting researchers to introduce novel applications Wang et al.
+(2020).
 For instance, Wang et al.
-(2022a) proposes long short-term extractors to quantify the win influence of each shot within a rally.
-Chang et al.
-(2022) designs the movement forecasting task to predict the players’ movements with graph-based approaches.
-In this paper, we aim to release, ShuttleSet22, a stroke-level badminton singles dataset collected from high-ranking matches in 2022.
-ShuttleSet22 is an extension from ShuttleSet Wang et al.
-(2023) encompassing with 30,172 strokes (2,888 rallies), 1,400 strokes (450 rallies) in the validation set, and 2,040 strokes (654 rallies) in the testing set.
-While ShuttleSet22 shares similar stroke-level data formats with ShuttleSet, ShuttleSet22 consists of matches in 2022 instead of between 2018 and 2021.
-Therefore, ShuttleSet can be viewed as the latest iteration of the badminton singles dataset that can be used to examine the model’s effectiveness in recent matches.
-ShuttleSet22 is obtained from the public videos [^2] http://bwf.tv/ and is labeled by domain experts with the shot-by-shot labeling tool Huang et al.
+(2022a) propose long short-term extractors to quantify the win influence of each shot within a rally, while Chang et al.
+(2023) design the movement forecasting task to predict players’ movements using graph-based approaches.
+In this paper, our aim is to introduce, ShuttleSet22, a stroke-level badminton singles dataset collected from real-world high-ranking matches in 2022.
+ShuttleSet22 extends the original ShuttleSet Wang et al.
+(2023), comprising 30,172 strokes (2,888 rallies), 1,400 strokes (450 rallies) in the validation set, and 2,040 strokes (654 rallies) in the testing set.
+While ShuttleSet22 shares similar stroke-level data formats with ShuttleSet, it consists of matches in 2022 instead of the period between 2018 and 2021.
+Therefore, ShuttleSet22 can be considered the most recent iteration of the badminton singles dataset, enabling the examination of model effectiveness in recent matches.
+ShuttleSet22 is sourced from public videos [^2] http://bwf.tv/ and has been meticulously labeled by domain experts with the shot-by-shot labeling tool Huang et al.
 (2022).
-To boost researchers’ engagements in badminton analytics, we, for the first time, create a challenge in CoachAI Badminton Challenge 2023333https://sites.google.com/view/coachai-challenge-2023/ in conjunction with IJCAI 2023.
-Specifically, we hold the forecasting of future turn-based strokes in badminton rallies (Track 2) and provide the state-of-the-art stroke forecasting approach Wang et al.
+Table 1: Comparison of the previous badminton dataset with our proposed ShuttleSet22.
+# Players # Match # Rally # Stroke Year
+BadmintonDB Ban et al. (2022) 2 9 811 9,671 2018-2020
+ShuttleSet22 (Ours) 35 58 3,992 33,612 2022
+To boost researchers’ engagements in badminton analytics, we have, for the first time, initiated a challenge within CoachAI Badminton Challenge 2023333https://sites.google.com/view/coachai-challenge-2023/ in conjunction with IJCAI 2023.
+Specifically, we have organized the forecasting of future turn-based strokes in badminton rallies (Track 2) and provided the state-of-the-art stroke forecasting approach Wang et al.
 (2022b) as the official baseline.
-This track has attracted about 100 participants to join for improving the effectiveness of the stroke forecasting task.
-The paper is organized as follows: Related work is described in Section 2, the details of the stroke forecasting task, the data exploration, and the official baseline are introduced in Section 3, methods of participants are summarized in Section 4, the results from participants are depicted in Section 5, and we conclude our findings in Section 6.
-We select it as the official baseline due to not only the state-of-the-art performance but also the first work for turn-based sequence prediction.
-2 Related Work: Public Badminton Datasets
-Table 1: Comparison of previous badminton datasets with our datasets.
-# Players # Match # Rally # Stroke Year
-ShuttleSet22 (Ours) 35 58 3,992 33,612 2022
-BadmintonDB Ban et al. (2022) 2 9 811 9,671 2018-2020
-Generally, there are only a few public badminton datasets due to the heavy cost of collecting and labeling fine-grained records by domain experts Wang (2022).
-Recently, researchers have released badminton datasets to foster the sports community.
-For instance, the shuttlecock datasets Cartron (2022) contains 8K images of shuttlecocks which are resized to 640*640 pixels.
-This dataset includes the position of shuttlecocks which could train an object detection model.
-The BadmintonDB Ban et al.
-(2022) features rally, strokes, and outcome annotations between two players, which can be used in player-specific match analysis and prediction tasks, which is also the stroke-level dataset.
-However, BadmintonDB only consists of the same matchup (i.e., Kento Momota and Anthony Sinisuka Ginting) instead of the various matchups.
-Moreover, BadmintonDB collects matches from 2018 to 2020, while our ShuttleSet22 collects high-ranking matches in 2022 to reflect the state-of-the-art tactic records.
-We summarize the discrepancies between our ShuttleSet and BadmintonDB in Table 1.
-3 Task Details
-Figure 1: Illustration of Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies.
-The goal of this shared task is to forecasting future turn-based strokes in badminton rallies, which aims to design forecasting models that are able to predict future strokes including shot types and locations based on the past stroke sequences.
-Figure 1 shows the whole flowchart of our task.
-The task page is available on Codalab.
-3.1 Problem Formulation
-Following Wang et al.
-(2022b), for each singles rally, given the observed $\tau$ strokes with type-area pairs and two players, the goal is to predict the future strokes including shot types and area coordinates for the next $n$ steps.
-To simplify the problem, $\tau$ is set to be 4 in our challenge, and $n$ is various based on the length of the rally.
-Let $R=\{S_{r},P_{r}\}_{r=1}^{|R|}$ denote historical rallies of badminton matches, where the $r$-th rally is composed of a stroke sequence with type-area pairs $S_{r}=(\langle s_{1},a_{1}\rangle,\cdots,\langle s_{|S_{r}|},a_{|S_{r}|}\rangle)$ and a player sequence $P_{r}=(p_{1},\cdots,p_{|S_{r}|})$.
-At the $i$-th stroke, $s_{i}$ represents the shot type, $a_{i}=\langle x_{i},y_{i}\rangle\in\mathbb{R}^{2}$ are the coordinates of the shuttle destinations, and $p_{i}$ is the player who hits the shuttle.
-We denote Player A as the served player and Player B as the other for each rally in this paper.
-For instance, given a singles rally between Player A and Player B, $P_{r}$ may become $(A,B,\cdots,A,B)$.
-We formulate the problem of stroke forecasting as follows.
-For each rally, given the observed $\tau$ strokes $(\langle s_{i},a_{i}\rangle)_{i=1}^{\tau}$ with players $(p_{i})_{i=1}^{\tau}$, the goal is to predict the future strokes including shot types and area coordinates for the next $n$ steps, i.e., $(\langle s_{i},a_{i}\rangle)_{i={\tau+1}}^{\tau+n}$.
-We note that $n$ is pre-defined as the actual length of the corresponding rally.
-3.2 Exploratory Data Analysis
-We had released 58 matches including about 4,000 rallies with shot-level records following the BLSR format Wang et al.
-(2022a) for this task, where parts of the matches from 2018 to 2021 are the same as Wang et al.
-(2022b), while we add new matches collected from 2022.
-Table 2: Explanation and example of each column.
-Column Name Explanation
-Rally Serial number of rallies in a match.
-Ball Round The order of the shot in a rally.
-Time The shot’s hitting time, which is in the format of hr:min:sec.
-Frame num Equals to sec multiply fps.
-Round score A Player A’s current score in the set.
-Round score B Player B’s current score in the set.
-Player The player who performed the shot.
-Type The type of shot.
-Aroundhead Hit the shuttle around the head or not.
-Backhand Hit the shuttle with the backhand or not.
-Landing height If the shuttle destination is hit above (1) or below (0) the net.
-Landing area The grid of the shuttle destinations
-Landing x
-Landing y
-The coordinates of the shuttle destinations.
-Reason for loss The reason why the rally ended.
-Get the point player The player who won the rally.
-Player location area The location of the player who performed the shot.
-Player location x,
-Player location y
-The coordinates of the player’s location when the player hits the shuttle.
-Opponent location area The location of the player who prepared to receive the shot.
-Opponent location x, Opponent location y The coordinates of the opponent’s location when the player hits the shuttle.
-Set The current set in a match (best-of-3).
-Match id Serial number of matches
-Rally id Serial number of all rallies
-Rally length Number of shots in a rally, rally_length-4 is the shots needed to predict.
-Table 2 summarizes the definition of each column, which is mainly followed by Wang et al.
-(2022a).
-Figure 2 shows the heatmap of player location and opponent location, which is obvious that most of the players stay in both centers of the court and four corners of the court when they are going to attack.
-On the contrary, most of the players stay in the center of the court when defending so that they can immediately react to any kind of shot.
-Table 3 indicates the number of different strike orders by shot type.
-It is observed that all shot types, except Clear and Smash, do not hit around the head.
-More than half of Push/Rush, Defensive Shot, and Short Service is hit with the backhand, while other shot types are hit with the forehand.
-The service type (i.e., how the player starts the first shot) has the most distinct difference in hitting position, with almost all short services hitting below the net and almost all long services hitting above the net.
-Figure 2: The heatmap of player location and opponent location in our datasets.
-Table 3: Shot type statistics.
+This track has attracted approximately 100 participants aiming to improve the effectiveness of the stroke forecasting task, with 16 teams submitting their final results on the leaderboard.
+In addition to the results, we summarize these insights and analyses of this challenge to inspire ideas for bridging the gap between badminton analytics and artificial intelligence communities [^4] The video can be found at https://youtu.be/yhRouMpxb2M..
+Table 2: The statistics of shot types. Short and long services only happen at the first stroke.
 Shot Type Not Around Head Around Head Forehand Backhand Higher than the net Below than the net Total Count
@@
 Long Service 648 0 528 120 644 3 648
+2 Related Work: Public Badminton Datasets
+Generally, there are only a few public badminton datasets due to the heavy cost of collecting and labeling fine-grained records by domain experts Wang (2022).
+Recently, researchers have released badminton datasets to foster the sports community.
+For instance, the shuttlecock datasets Cartron (2022) contains 8K images of shuttlecocks which are resized to 640*640 pixels.
+This dataset includes the position of shuttlecocks which could train an object detection model.
+The BadmintonDB Ban et al.
+(2022) features rally, strokes, and outcome annotations between two players, which can be used in player-specific match analysis and prediction tasks, which is also the stroke-level dataset.
+However, BadmintonDB only consists of the same matchup (i.e., Kento Momota and Anthony Sinisuka Ginting) instead of the various matchups.
+Moreover, BadmintonDB collects matches from 2018 to 2020, while our ShuttleSet22 collects high-ranking matches in 2022 to reflect the state-of-the-art tactic records.
+We summarize the discrepancies between our ShuttleSet and BadmintonDB in Table 1.
+3 The Challenge System
+Figure 1: Illustrated system of Track 2: Forecasting Future Turn-Based Strokes in Badminton Rallies.
+Conventional applications in badminton mainly focus on quantifying stroke performance Wang et al.
+(2021) or retrieving information from videos Chu and Situmeang (2017), which motivates us to further investigate a more challenging yet critical real-world application.
+The goal of this shared task is to forecasting future turn-based strokes in badminton rallies, which aims to design forecasting models capable of predicting future strokes, including shot types and locations, based on past stroke sequences.
+Figure 1 illustrates the entire flowchart of our task.
+As we have built the pipeline from data processing to evaluations, participants only need to modify the stroke forecasting model to swiftly iterate their approaches.
+The task page is available on Codalab.
+Figure 2: Heatmap of player and opponent locations in ShuttleSet22.
+3.1 Problem Formulation
+Following the definition outlined in Wang et al.
+(2022b), for each singles rally, given the observed $\tau$ strokes with type-area pairs and two players, the goal is to predict the future strokes, including shot types and area coordinates, for the next $n$ steps.
+To simplify the problem, $\tau$ is set to be 4 in our challenge, and $n$ is various based on the length of the rally, which is given to the participants.
+Let $R=\{S_{r},P_{r}\}_{r=1}^{|R|}$ denote historical rallies of badminton matches, where the $r$-th rally is composed of a stroke sequence with type-area pairs $S_{r}=(\langle s_{1},a_{1}\rangle,\cdots,\langle s_{|S_{r}|},a_{|S_{r}|}\rangle)$ and a player sequence $P_{r}=(p_{1},\cdots,p_{|S_{r}|})$.
+At the $i$-th stroke, $s_{i}$ represents the shot type, $a_{i}=\langle x_{i},y_{i}\rangle\in\mathbb{R}^{2}$ are the coordinates of the shuttle destinations, and $p_{i}$ is the player who hits the shuttle.
+We denote Player A as the served player and Player B as the other for each rally in this paper.
+For instance, given a singles rally between Player A and Player B, $P_{r}$ may become $(A,B,\cdots,A,B)$.
+We formulate the problem of stroke forecasting as follows.
+For each rally, given the observed $\tau$ strokes $(\langle s_{i},a_{i}\rangle)_{i=1}^{\tau}$ with players $(p_{i})_{i=1}^{\tau}$, the goal is to predict the future strokes including shot types and area coordinates for the next $n$ steps, i.e., $(\langle s_{i},a_{i}\rangle)_{i={\tau+1}}^{\tau+n}$.
+We note that $n$ is pre-defined as the actual length of the corresponding rally.
+3.2 Exploratory Data Analysis
+We have released 58 matches, approximately 4,000 rallies with shot-level records following the BLSR format Wang et al.
+(2022a) for this task, where parts of the matches from 2018 to 2021 are identical to those in Wang et al.
+(2022b), while we have further included new matches collected from 2022.
+Figure 2 shows the heatmap of player location and opponent location, indicating that most players position themselves at both the center and the four corners of the courts when preparing to attack.
+On the contrary, when defending, players intend to stay predominantly in the center of the court to enable quick reactions to any type of shot.
+Table 2 presents the numbers of different strike orders by shot types.
+It is evident that with the exception of the Clear and Smash shots, other shot types do not target around the head area.
+Furthermore, more than half of the Push/Rush, Defensive Shot, and Short Service shots are executed with the backhand, while other shot types predominantly utilize the forehand.
+Notably, there is a significant disparity in hitting position based on the service type, with almost all short services directed below the net and almost all long services aimed above the net.
 3.3 Evaluation Metrics
-The evaluation scenarios will be evaluated by cross-entropy for shot type prediction and mean absolute error (MAE) for area coordinates prediction similar to the original work.
-Since this task exists stochasticity, we will also ask each team to generate 6 predicted sequences of each rally and take the closest one to ground truth for evaluation.
-We note that the original work of the stroke forecasting task is to generate 10 sequences, which we change to 6 for efficiency.
+The evaluation scenarios will be assessed by cross-entropy for shot type prediction and mean absolute error (MAE) for area coordinates prediction, similar to the original work.
+Given the stochastic nature of this task, each team will be required to generate 6 predicted sequences for each rally, from which the closest one to the ground truth will be selected for evaluation.
+It is worth noting that the original work on the stroke forecasting task involved generating 10 sequences, a number we have reduced to 6 for efficiency.
 3.4 Official Baseline
@@
 (2022b) is a position-aware fusion of rally progress and player styles framework consisting of Transformer-based architectures.
-ShuttleNet separates the contexts of both players to consider the overall playing habits for predicting the next strokes.
-4 Participating Systems
-There are about 95 participants joining the challenge system, and 16 teams submitted their testing results for the final phase in the CoachAI Badminton Challenge 2023 (Track 2).
-Most teams modified the official baseline (i.e., ShuttleNet) to approach the task.
-The brief descriptions of the submitted methods are summarized in Table 4.
-5 Results
-Table 4: Performance and brief descriptions of the stroke forecasting task in CoachAI Badminton Challenge 2023 (Track 2).
-$\tau=4$
-Rank Team Descriptions CE MAE Total
-1 Intro_to_AI_team8 Since the embedded player number appears in the original ShuttleNet structure multiple times, we doubt that we can remove some of them to make our model simpler.
-Similarly, the embedded area has gone through many sigmoid and relu functions repetitively before passing them to the neural network.
-We simply ignore all the unnecessary activation functions and embedding processes in the original structure.
-1.7892 0.7884 2.5776
-2 Badminseok The given input parameters include ”type,” ”aroundhead,” ”backhand,” ”landing height,” ”landing_x,” ”landing_y,” ”player_location_area,” and ”opponent_location_area.” The desired output parameters are the same as the input parameters.
-A total loss function is defined as a weighted combination of two components: ”loss_area” and ”loss_type.” The weights are controlled by the alpha value, which can be adjusted for tuning.
-Additional losses, such as ”loss_aroundhead,” ”loss_backhand,” ”loss_height,” ”loss_playerlocationarea,” and ”loss_opponentlocationarea,” are included.
-The final ideal loss is obtained by combining the weighted losses of the two models.
-Hyperparameter tuning is performed separately for model 1 (optimizing loss area) and model 2 (optimizing loss type).
-The ensemble model is created by combining the outputs of both models to generate the final data.
-Additionally, we introduced two forward layers that returns vectors of area and shot to deepen the encoder and the decoder depth.
-The modified model showed better performance on regression task.
-1.8127 0.7703 2.5830
-3 NYCU-group4 Optimizing the parameters of the original baseline code, and testing the accuracy of the model on validation sets to obtain the model with the optimal weights. 1.8411 0.7826 2.6237
-4 YHY We basically use the ShuttleNet from the paper to predict the future steps of type and landing coordinate. 1.9685 0.6797 2.6482
-5 20 Use 9 convolution in FFN and add relative positional bias. 1.9366 0.7490 2.6856
-6 LOL data augmentation using spatial mirroring and finding matching technique, added player area as input feature, change of activation function from baseline code. 1.9390 0.7618 2.7008
-7 Group27 Based on ShuttleNet, we add Convolutional Unit and use Bayesian Optimization to find the optimal hyperparameter. 2.0200 0.7139 2.7339
-8 Team_13 Based on Shuttlenet and adjust parameters. 1.9681 0.7671 2.7352
-9 LinDan Train on extra epochs. 2.0097 0.7743 2.7841
-10 Intro_to_AI_group_5 We apply some feature extraction methods to integrate features provided in the dataset.
-In comparison to the original ShuttleNet model, our work generates predictions based on more features, including shot angle, time difference, etc.
-1.9710 0.8726 2.8436
-11 14 We use ShuttleNet as our main approach in this competition. 2.1718 0.7013 2.8731
+ShuttleNet has demonstrated superior performance in predicting the next strokes compared to conventional sequential.
+This is attributed to its turn-based architecture, which separates the styles of both players in a rally and integrates them with the current rally condition.
+Table 3: Performance of the stroke forecasting task in CoachAI Badminton Challenge 2023 (Track 2).
+Rank Team CE MAE Total
+1 Intro_to_AI_team8 1.7892 0.7884 2.5776
+2 Badminseok 1.8127 0.7703 2.5830
+3 NYCU-group4 1.8411 0.7826 2.6237
+4 YHY 1.9685 0.6797 2.6482
+5 20 1.9366 0.7490 2.6856
+6 LOL 1.9390 0.7618 2.7008
+7 AI Project #15 1.9536 0.7483 2.7019
+8 Group27 2.0200 0.7139 2.7339
+9 Team_13 1.9681 0.7671 2.7352
+10 LinDan 2.0097 0.7743 2.7841
+11 Intro_to_AI_group_5 1.9710 0.8726 2.8436
+12 14 2.1718 0.7013 2.8731
 - ShuttleNet Wang et al. (2022b) 2.1777 0.6997 2.8774
-12 GD_Wang Discretize location for ball type attention & curriculum learning with multi-task (base on ShuttleNet). 2.2479 0.7081 2.9560
-13 ShuttleFold We commence our methodology by detailing the dataset employed and the associated data preprocessing steps.
-Following this, drawing upon the resemblances with AlphaFold2, we elucidate the initialization of our unique player and rally representations.
-Subsequently, we delve into the intricacies of the Evoformer, which facilitates the update of these representations.
-2.1277 1.0107 3.1385
-14 Awesome Badminton Increase embeddings dimension. 2.6579 0.7120 3.3699
-15 Badminton is all you need Autoregressive model with SOTA techniques that will reveal soon, only trained in under 200 epochs. 2.6579 0.7120 3.3699
-16 Group_28 Select several significant features for the past 4 stroke sequence and train a model with the encoder-decoder architecture for a single future sequence prediction including landing coordinates.
-Use the selected- predicted features as attributes for NN categorical classification on type of strokes.
-We iterates over the two continuously until the ball round ends.
-4.6615 1.0268 5.6883
+13 GD_Wang 2.2479 0.7081 2.9560
+14 ShuttleFold 2.1277 1.0107 3.1385
+15 Awesome Badminton 2.6579 0.7120 3.3699
+16 Badminton is all you need 2.6579 0.7120 3.3699
+17 Group_28 4.6615 1.0268 5.6883
+4 Participating System
+Approximately 95 participants have joined the challenge system, with 16 teams submitting their testing results for the final phase of the CoachAI Badminton Challenge 2023 (Track 2).
+Most teams modified the official baseline, ShuttleNet, to address the task.
+The brief descriptions of the submitted methods are summarized in the video due to the page limit.
+5 Results and Findings
 Table 3 reports the official leaderboard of the participants’ methods.
-We can observe that Team Intro_to_AI_team 8 outperforms the competition slightly than Team Badminseok, while Team YHY performs the best in terms of area predictions.
-There are 11 teams that perform better than ShuttleNet; however, we notice that these methods are built on top of ShuttleNet, which demonstrates the flexibility of ShuttleNet and some designs (e.g., hyper-parameters and activation functions) can be improved.
-In addition, the improvements of the participating teams mainly focus on the shot type prediction (2.1777 $\rightarrow$ 1.7892); however, the performance of area coordinates only slightly performs better than ShuttleNet (0.6997 $\rightarrow$ 0.6797), where most teams are inferior to the baseline.
-This points out the challenge of effectively incorporating two predictions, which is a potential future direction to explore.
+It is observed that Team Intro_to_AI_team 8 slightly outperforms the competition Team Badminseok, while Team YHY demonstrates the best performance in terms of area predictions.
+11 teams perform better than ShuttleNet; however, we notice that these methods are built on top of ShuttleNet, which highlights the flexibility of ShuttleNet and the potential for improvements in certain aspects, e.g., hyper-parameters and activation functions.
+In addition, the improvements made by the participating teams primarily focus on the shot type prediction (2.1777 $\rightarrow$ 1.7892); however, the performance of area coordinates only marginally surpasses ShuttleNet (0.6997 $\rightarrow$ 0.6797), with most teams being inferior to the baseline.
+This underscores the challenge of effectively integrating two predictions, which may be a potential future direction for exploration.
 6 Conclusion
 In this paper, we propose ShuttleSet22, an extended dataset from ShuttleSet with stroke-level badminton singles records.
-ShuttleSet22 consists of fine-grained metadata for researchers to utilize various aspects for discovering thoughtful insights.
-Exploratory data analysis is conducted to illustrate the fundamental compositions of the dataset to bridge the gap between researchers outside the badminton domain.
-To foster researchers to incorporate advanced techniques with badminton analytics, we hold Track 2, a challenge in CoachAI Badminton Challenge 2023 that focuses on improving the effectiveness of stroke forecasting.
-In addition, we replicate the state-of-the-art baseline for the task as the official baseline.
-Methods and insights from participants are summarized and discussed to provide potential improvements for future research.
+ShuttleSet22 consists of fine-grained metadata to reinforce researchers to explore various aspects and derive insightful findings.
+Exploratory data analysis is conducted to elucidate the fundamental compositions of the dataset, aiming to bridge the gap between researchers outside the badminton domain.
+To foster researchers to incorporate advanced techniques into badminton analytics, we introduce Track 2, a challenge within CoachAI Badminton Challenge 2023, focused on enhancing the effectiveness of stroke forecasting.
+In addition, we establish the state-of-the-art baseline for the task.
+We summarize and discuss the methods and insights from participants to provide potential avenues for future research improvement.